Farrelly Enterprises has fixed operating costs of $500,000, variable operating costs per unit of $20, and a selling price of $40 per unit. It produces 40,000 units. Company has a $30,000 bond with a 10% coupon rate and it has issued 400 preferred shares with a $3 per share dividend payment. Farrelly also has 1000 outstanding shares of common stock and it falls in the 34% tax bracket.
a. Prepare an income statement showing the percentage change in EPS if sales increase by 10%. Calculate the Degree of Total Leverage using the percentage change formula.
b. Calculate the Degree of Operating Leverage and Degree of Financial Leverage using the direct formula.
1
Expert's answer
2020-09-06T17:28:42-0400

a.


"Degree of Total Leverage=DOL\times DFL"

DOL is the degree of operating leverage and DFL is the degree of financial leverage.

It can be established that DOL equals contribution margin (i.e. sales minus variable costs) divided by earnings before interest and taxes (EBIT) and DFL equals EBIT divided by (EBIT minus interest expense I).

"DOL=\frac{1 760 000-800 000}{460 000}=\frac{960 000}{460 000}=2.09"


"DFL=\frac{460 000}{460 000-3000}=\frac{460 000}{457000}=1.01"


"Degree of Total Leverage=DOL\times DFL=2.09\times1.01=2.11"

b."Degree of Operating Leverage =\frac{ (Sales – Variable cost)}{(Sales – Fixed cost – Variable cost)}=\frac{1760 000-800 000}{1760 000-500 000-800 000}=\frac{960 000}{460 000}=2.09"


"Degree of Financial Leverage =\frac{Sales – Fixed cost – Variable cost}{Sales – Fixed cost – Variable cost-interests bonds}=\frac{1760 000-500 000-800000}{1760 000-500 000-800 000-3000}=\frac{460 000}{457 000}=2.11"